### v1

# A Large Scale Analysis of Gender Biases in Text-to-Image Generative Models

Abstract
With the increasing use of image generation technology, understanding its social biases, including gender bias, is essential. This paper presents the first large-scale study on gender bias in text-to-image (T2I) models, focusing on everyday situations. While previous research has examined biases in occupations, we extend this analysis to gender associations in daily activities, objects, and contexts. We create a dataset of 3,217 gender-neutral prompts and generate 200 images per prompt from five leading T2I models. We automatically detect the perceived gender of people in the generated images and filter out images with no person or multiple people of different genders, leaving 2,293,295 images. To enable a broad analysis of gender bias in T2I models, we group prompts into semantically similar concepts and calculate the proportion of male- and female-gendered images for each prompt. Our analysis shows that T2I models reinforce traditional gender roles, reflect common gender stereotypes in household roles, and underrepresent women in financial-related activities. Women are predominantly portrayed in care- and human-centered scenarios, and men in technical or physical labor scenarios. Code and data will be released.
Figure 1: Visualization of our image generation setup to analyze gender bias in T2I models with respect to everyday scenarios.
1 Introduction
Rapid advances in image generation technology make it easier than ever to automatically generate large amounts of synthetic images. State-of-the-art text-to-image (T2I) models [52, 83] can generate high-quality images from arbitrary text instructions. Their capabilities are further enhanced through editing [10, 49, 101] and personalization [75, 34, 50, 9] techniques. Synthetic images are not only used in everyday applications such as advertisements [53] and presentation slides [67] but are also increasingly used as training data for other foundation models [86, 87, 30, 50].
As the availability and proliferation of synthetic images increase, so does their power to influence society and amplify any harms originating from the underlying models [15]. In their seminal work, [12] discovered intersectional gender and racial biases in image recognition systems. Within the research community, the list of known biases has only grown: [5, 38, 7, 81, 84] identified social biases in CLIP [69], such as associating men with words related to criminal activities. [40, 45, 46] found social biases in automatic image captioning. [100, 96, 36, 32, 74] uncovered various social biases in multimodal large language models (MLLMs). These examples demonstrate that social bias pervades all aspects of modern generative AI systems.
Research on social bias in T2I models has led to a large body of work covering all computational aspects of social bias, including bias analysis [21, 8, 56], open bias detection [24, 20, 23], and model debiasing [99, 6, 28, 33]. However, most research in this area has focused on gender-occupation bias [89]. While this is an important issue, other aspects of daily life, such as everyday activities and stereotypical contexts, also contribute to perpetuating or amplifying harmful social biases and require careful analysis. Furthermore, many studies on social bias in T2I models use a limited set of prompts and only a few images per prompt, reducing their representativeness.
This work presents the first large-scale study on gender bias in T2I models concerning everyday contexts. We address gaps in the literature by analyzing everyday activities complemented by gender-object and gender-context associations. Our main contributions are: (1) We compile 3,217 gender-neutral prompts over four categories to probe T2I models for gender bias as illustrated in Fig. 1; (2) We generate 200 images from five state-of-the-art T2I models for each prompt and filter unsuitable images, leaving 2,293,295 images for analysis; (3) We design a carefully structured experimental setup to analyze gender bias in large image datasets as detailed in Fig. 2; (4) We systematically examine the observed gender biases and relate them to known human gender biases. Through this work, we aim to take an important step toward addressing gender stereotypes in T2I models, thereby helping to understand perpetuation of gender inequality in generative AI technology [85, 2].
2 Related Work
Prior work on evaluating social bias in text-to-image models can be categorized into bias analysis and open bias detection. Notable works on open bias detection include TIBET [20], OpenBias [24], and OASIS [23]. In this work, we conduct a careful, large-scale study of gender bias in state-of-the-art diffusion models. This paper focuses specifically on analyzing gender bias in text-to-image models by exploring gendered outputs in gender-neutral prompts.
Work analyzing social bias in text-to-image (T2I) generation has traditionally focused on a few categories [89], most notably perceived gender and race, which are protected against discrimination under U.S. federal law. In their seminal work, [8] find that T2I models amplify bias across many categories. [19] examine gender and racial bias in occupations and report that generated images are more gender-imbalanced than actual U.S. labor statistics. However, [80] provides evidence that bias amplification may stem from specific prompt design choices. [56] propose a method to analyze gender and racial bias without explicitly labeling images by gender or race, which can be controversial due to assigning gender and race to images. They find distributional differences in images generated for occupation-related prompts. Similarly, [21] observe distributional differences when generating images of occupations using gendered and gender-neutral prompts. Other studies explore different aspects of bias. [102, 93, 59] investigate gender bias in person-object co-occurrence. [88] find that non-binary identities are poorly represented in T2I models. Likewise, [35] show that T2I models inadequately represent national identities and tend to produce overly sexualized images of women, particularly when prompted for regions in the Global South. Expanding this perspective, [48] quantifies national stereotyping in generated images.
Previous work has established that T2I models exhibit gender bias, among other issues. However, as also pointed out by [89], scenarios beyond occupation are underrepresented in current research. Furthermore, most studies are based on small-scale analyses, typically using fewer than 200 prompts and no more than 20 images per prompt. Benchmarks that aim for a larger scale, such as [57], are still dominated by occupation-related prompts. Thus, this work provides the means for large-scale analyses of gender bias beyond occupations. Additionally, [58] find that many images generated from commonly used prompt templates do not depict people with distinguishable gender characteristics, so we ensure the validity of our analysis by filtering images without people or showing both men and women.
Figure 2: Overview of our experimental setup to analyze gender bias in T2I models regarding everyday scenarios. We show our 4 prompt groups on the bottom right and the five T2I models on the bottom left. The top left visualizes our filtering method: First, we detect people, crop the bounding boxes, and detect perceived gender. We remove images without people or showing at least one man and one woman. We calculate the proportions of female- and male-labeled images generated for each prompt and analyze systematic gender biases.
3 Prompts and Images
3.1 Prompt Collection, Processing and Clustering
We collect prompts in four categories: (1) Activities, (2) Contexts, (3) Objects, and (4) Occupations. To study gender bias in everyday activities, we rely on the curated set from [92], i.e. Activities. The activities in [92] were gathered through Amazon Mechanical Turk from U.S. based workers, who were asked to provide short phrases describing recent activities. Including these 1405 activities provides insight into how stereotypical gender roles in everyday life are reflected in T2I models. Analyzing gender associations with contexts and objects further extends this analysis by considering everyday situations beyond specific activities. To study gender bias in relation to people and places, i.e. Contexts, we collect a set of 737 scene classes from the SUN Database [94, 95]. We include all classes but do not consider fine-grained distinctions, e.g. “inside” the church and “outside” the church leads to the context “church”.
We collect 500 common physical objects from WordNet [31], i.e. Objects. To select these 500 objects, we filter all noun hyponyms of the object.n.01 WordNet synset using a list of the most common English words. Additionally, we manually remove a small number of unsuitable synsets, e.g. synsets that refer to people or body parts, or places that are already covered in the contexts prompt group. We retain the top 1000 most frequent lemmas and re-rank them based on their concreteness following [11]. Finally, we select the top 500 most concrete lemmas from the top 1000 most frequent WordNet lemmas. We also include occupations to align with prior work and to examine occupation-related gender bias in T2I models. Our occupation list is comparatively large, as we include all 575 occupations listed by the U.S. BLS [64] rather than a subset, i.e. Occupations.
Using an LLM, specifically Yi-1.5-34B [98] (see Sec. A.2 for a comparison to other LLMs), we process all collected activities, contexts, objects, and occupations to generate syntactically coherent prompts. Since we do not specify gender in our prompts, each prompt begins with “a person”. We then apply a different template for each prompt group as shown in Tab. 1. Detailed prompts for filling in the templates can be found in Sec. A.1. Additionally, we simplify occupation descriptions, which are often overly detailed in [64].
Activity a person who is {{verb}}-ing {{activity}}”
Context a person {{in/on/at…}} the {{context}}”
Objects a person and {{a object}}”
Occupations a person working as {{occupation}}”
Table 1: Prompt templates for the different prompt groups. Parts in double brackets {{}} are modified or filled in by the LLM.
We generate 5 variations per prompt by replacing the prefix “a person” with “an individual”, “someone”, “a friend”, and “a colleague”, which do not include any gender information. Prompt variations increase diversity and are essential for valid analysis of social bias in large models [79, 78, 42]. Sec. B.3 shows that variations do not lead to a significant skew towards male- or female-gendered images.
Figure 3: Stacked distribution of female ratios in generated images for all models and prompt groups.
For a concise presentation of our findings from the 3217 prompts, we cluster prompts in all four prompt groups into semantically coherent concepts to facilitate analysis. We apply the following variation of BERTopic [37] to cluster prompts. First, we embed all prompts using a sentence embedding model [70]. Then, we reduce the embedding dimensions to 16 using UMAP [62]. Using HDBSCAN clustering [61] with cosine distance, we determine the concept clusters. However, HDBSCAN has the option to not assign a prompt to any cluster, so we add unclustered prompts to the cluster with nearest centroid. The detailed settings are in Sec. A.3. We arrive at 165 activity clusters, 91 context clusters, 62 object clusters, and 76 occupation clusters.
After clustering prompts, we summarize all clusters by an LLM (see Sec. A.3). Summarizing a list of prompts requires more in-depth understanding and reasoning than prompt processing, so we use Llama-3-3-70B-Instruct based on manual inspection. A comparison to other LLMs and our exact prompt template is in Sec. A.3. For the purposes of our analysis, we further merge clusters with the same summary (e.g. different variants of shopping-related activities), as they would be indistinguishable for the reader. However, we will make the fine-grained clusters available alongside our code.
3.2 Image Generation and Gender Identification
We generate images using 5 models, which represent the state-of-the-art among open models at the time of writing: (1) Flux [52], (2) Flux-Schnell, (3) Stable Diffusion 3.5 Large [83], (4) Stable Diffusion 3.5 Medium, and (5) Stable Diffusion 3 Medium [29]. These are latent diffusion models [72] based on Diffusion Transformers [66]. Other recent strong models, such as Lumina 2.0 [55] and Janus-Pro-7B [16], were concurrently released to this study and unavailable when conducting experiments. For each combination of model and prompt, we generate 40 images per prompt variation. This results in 5 models $\times$ 5 prompt variations $\times$ 3217 prompts $\times$ 40 images = 3,217,000 images.
To analyze gender bias in generated images, we identify the perceived gender of the people shown in the images using a two-step process. First, we detect all people in the generated images using the object detection model YOLOv10 [90] and obtain bounding boxes for the detected individuals. Next, we crop each detected person’s bounding box and pass it to an MLLM, InternVL2-8B [17, 18], along with a prompt asking the model to identify the person’s gender as “female”, “male”, or “unclear/cannot tell”. We focus on binary perceived gender, specifically men and women, for several reasons. First, current T2I models do not generate distinct features that clearly indicate non-binary gender. Second, current MLLMs rarely consider non-binary gender as an option, as shown in Sec. C.2. While not addressed in this paper, the fact that models output gender as binary is a separate issue that warrants further discussion.
Using bounding boxes instead of the full image helps mitigate bias from the person’s context, i.e. predicting the gender based on the background and not the person’s features, as MLLMs also exhibit gender bias [36]. It also prevents confusion when multiple people of potentially different genders appear in the image. In Sec. C.1, we provide the detailed prompt used for gender identification and verify, using human-labeled data, that the MLLM used in this study can identify perceived gender with near-perfect accuracy. It is important to note that assigning a person’s gender can be problematic, because it cannot necessarily be perceived from an image, and gender is a spectrum. Therefore, good practice with images of real people is to have people self-identify their gender. However, T2I models create images that are not real, so assigning perceived gender to the images is more acceptable as there is no risk of misidentifying a real person.
Activities Contexts Objects Occupations
# Prompts (# Imgs.) 1405 (1,405K) 737 (737K) 500 (500K) 575 (575K)
Flux 1149 (187,079) 632 (100,125) 395 (58,387) 574 (110,989)
Flux-Schnell 1096 (168,994) 693 (105,737) 466 (63,520) 574 (105,288)
SD-3.5-large 1163 (185,310) 689 (113,839) 476 (77,789) 570 (108,454)
SD-3.5-Medium 1076 (163,845) 693 (112,296) 411 (59,507) 572 (107,691)
SD-3-Medium 1062 (169,403) 711 (124,520) 418 (62,702) 573 (107,820)
Table 2: Prompt Groups with remaining prompts after filtering, and the exact number of remaining images in brackets.
3.3 Image and Prompt Filtering
We exclude images and prompts that are not suitable for analyzing gender bias, i.e. if (a) There is no person in the image; (b) There is no person in the image whose gender can be clearly identified (see Sec. C.3 for details); (c) There are multiple people in the image and there is at least one man and one woman. If there are multiple people, we keep the image if people with all the same gender are shown or where the gender of other people is labeled “unclear/cannot tell.” (for example people in the background).
Additionally, we exclude entire prompts for a model if fewer than 100 out of 200 images remain after filtering. Since we analyze gender bias at the prompt level, we can only consider prompts where we can reliably estimate the ratio of male and female people in the images. The number of remaining prompts for each model is in Tab. 2.
Figure 4: (Top) Top 10 most female-dominated (top row) and top 10 most male-dominated (bottom row) activity clusters. Bars indicate the ratio of female-gendered images generated from 5 T2I models averaged over prompts in each cluster, and the error line indicates the standard deviation across prompts. (Bottom) Top 5 most female-dominated (left) and top 5 most male-dominated (right) context clusters.
4 Gender Bias Analysis Experiments
For each prompt, e.g. the Activities prompt group has 1405 prompts, we calculate the ratio of male and female images. Let $\mathcal{I}^{p}=\{I_{1}^{p},I_{2}^{p},\ldots,I_{n}^{p}\}$, $n\leq 200$ be the set of gendered images for prompt $p$ after filtering and $\mathcal{G}:\mathcal{I}\rightarrow\{\textrm{female},\textrm{male}\}$ the mapping of images to unique genders according to InternVL2-8B. Then, we define the female-gendered images $F(p)$ and male-gendered images $M(p)$ as
$F(p)$ :=$:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{female}\}$ (1)
$M(p)$ :=$:=\{I\in\mathcal{I}^{p}\ |\ \mathcal{G}(I)=\textrm{male}\}$ (2)
and the female ratio $\mathcal{R}_{f}(p)$ of prompt $p$ as
$\mathcal{R}_{f}(p):=\frac{\left\lvert F(p)\right\rvert}{\left\lvert F(p)\right\rvert+\left\lvert M(p)\right\rvert}.$ (3)
This allows us to estimate the distribution of female ratios across activities for a given model, i.e. we present the distribution of values of $\mathcal{R}_{f}$ as a histogram in Fig. 3. Overall, we find that the models generate similar gender ratios across all prompts (see Appendix D for more details). However, we also observe that models tend to generate more male-gendered images, as also observed by [35, 99].
4.1 Activities and Contexts
In Fig. 3, we find a large number of activities in which the set of images is male-dominated with $\mathcal{R}_{f}$ close to zero. We say a prompt or a prompt cluster is female-dominated (male-dominated) if $\mathcal{R}_{f}\geq 0.7$ ($\mathcal{R}_{f}\leq 0.3$), i.e. 70% (30%) or more (less) of images for this prompt or cluster are female-gendered. If $\mathcal{R}_{f}$ is between 50% and 70%, we speak of female-leaning clusters or prompts (equivalently male-leaning). While gender ratios of activities are distributed more evenly, the distribution of gender ratios for contexts is heavily skewed toward male images. This highlights a general trend to outputting males overall and men as the default. We calculate the top 10 (top 5) activities with the highest ratio of female- or male-gendered images across T2I models to showcase male- and female-dominated activities (contexts). We state the summary and the per-model average ratio of female- or male-gendered images of activities (contexts) in the given cluster for each activity (context) cluster. Results are in Fig. 4. For each cluster, we also report the average value of $\mathcal{R}_{f}$ across models (^$\hat{\mathcal{R}}_{f}$).
Most female-dominated activities. Common female-dominated activities are crafting (^$\hat{\mathcal{R}}_{f}\approx 85\%$), which comprises activities such as “crocheting” or “making bracelets”, and pet-related activities (cat (^$\hat{\mathcal{R}}_{f}\approx 79\%$), pet (^$\hat{\mathcal{R}}_{f}\approx 81\%$)). birthday (^$\hat{\mathcal{R}}_{f}\approx 83\%$) contains activities related to parties. Further care-related activities are baby care (^$\hat{\mathcal{R}}_{f}\approx 88\%$) and volunteer (^$\hat{\mathcal{R}}_{f}\approx 82\%$), which are both examples of helping other people. The prominence on care-related activities reflects gender norms of women as caretakers and resembles human stereotypes, as women are described as “warm”, “sensitive to others”, and specifically “interested in children” [73]. The remaining highly female-dominated clusters in Fig. 4 are shopping-related activities (shopping (^$\hat{\mathcal{R}}_{f}\approx 83\%$), grocery (^$\hat{\mathcal{R}}_{f}\approx 76\%$)) and yoga-related activities in physical (^$\hat{\mathcal{R}}_{f}\approx 83\%$). Shopping-related activities include both shopping for daily necessities and clothes. Grocery shopping is known to be a household activity typically performed by women [22], and clothes are associated with women also in other prompt groups, especially contexts and objects (Sec. 4.2). Yoga was found to be seen as strongly female-typed [60]. Finally, “baking” (^$\hat{\mathcal{R}}_{f}\approx 76\%$) is a female-typed way of cooking [71].
Most male-dominated activities. One common male-dominated activity type in Fig. 4 is outdoor household work (mowing (^$\hat{\mathcal{R}}_{f}\approx 3\%$), yard work (^$\hat{\mathcal{R}}_{f}\approx 16\%$)), which includes “mowing the lawn”, “cutting wood”, and “raking leaves”. Mowing the lawn specifically was identified as an activity typically performed by men [22]. Further male-dominated activities are car-related (car, car maintenance (both ^$\hat{\mathcal{R}}_{f}\approx 8\%$)), which is also male-typed [22], as well as media consumption, such as (computer) gaming (^$\hat{\mathcal{R}}_{f}\approx 9\%$) and movie watching (^$\hat{\mathcal{R}}_{f}\approx 10\%$) or tv viewing (^$\hat{\mathcal{R}}_{f}\approx 13\%$). According to [43], young men, on average, devote more time to “watching TV and video” and “computer games” than women of the same age. However, [82, 65] find that (computer) gamers being predominantly male is more of a stereotype than reality, meaning that T2I models perpetuate the marginalization of women in e-sports [65]. Smoking (^$\hat{\mathcal{R}}_{f}\approx 8\%$) explicitly refers to cannabis consumption, which, alongside other drug consumption including alcohol, is more common among men than women [91, 77, 39]. Football (^$\hat{\mathcal{R}}_{f}\approx 2\%$) is a strongly male-typed sport [68] and also strongly male-dominated in T2I models. Many male-gendered images in the parent-son (^$\hat{\mathcal{R}}_{f}\approx 2\%$) cluster are less surprising as prompts contain gendered words, i.e. “son”.
Most female-dominated contexts. The only consistently female-leaning clusters are school (^$\hat{\mathcal{R}}_{f}\approx 68\%$) and student living (^$\hat{\mathcal{R}}_{f}\approx 55\%$), describing university environments, such as “classroom”, as well as childcare (^$\hat{\mathcal{R}}_{f}\approx 60\%$), containing places such as “playroom”. Teachers is a profession with female majority [64], which could explain the association of women with school places. Retail (^$\hat{\mathcal{R}}_{f}\approx 55\%$) contains various shopping locations, of which only a subset is strongly female-dominated. Such locations are related to fashion, such as “jewelry shop” or “hat shop”. Details on the retail cluster are in Sec. E.1.
Most male-dominated contexts. In contrast to female-dominated contexts, the most male-dominated clusters focus on transportation (auto (^$\hat{\mathcal{R}}_{f}\approx 7\%$), shipping (^$\hat{\mathcal{R}}_{f}\approx 10\%$), parking and storage (^$\hat{\mathcal{R}}_{f}\approx 10\%$)) and industrial places (work (^$\hat{\mathcal{R}}_{f}\approx 9\%$)) Another strongly male-dominated cluster is baseball (^$\hat{\mathcal{R}}_{f}\approx 7\%$), which contains “baseball field” and various locations therein, such as “batting cage” or “pitcher’s mound”. We note that baseball is a male-leaning sport [60]. It is clear that T2I models do not associate women with industrial, work-related places, and places where women are depicted seem to be social places, such as schools or certain shops. In Sec. F.2, we provide further analyses of workplace gender bias.
Figure 5: (Top) Top 5 most female-dominated (left) and top 5 most male-dominated (right) object clusters. (Bottom) Top 5 most female-dominated (left) and top 5 most male-dominated (right) occupation clusters (note that here the y-axis is between 0% and 10%).
4.2 Objects and Occupations
In Fig. 3, we observe a roughly Gaussian distribution for objects, peaking at a 0.4 female ratio. There are relatively few objects where models generate exclusively male- or female-gendered images. Gender distributions for occupations are highly polarized, with most occupations yielding only male-gendered images. Compared to actual work participation statistics, this reflects a clear bias amplification, in line with the findings in [80]. However, it also means that most previous work focusing on occupations has studied a particularly extreme example of bias amplification in T2I models. This further justifies our focus on activities and other everyday contexts. We list the top 5 object (occupation) clusters by female- and male ratios for each T2I model in Fig. 5.
Most female-dominated objects. Main theme in female-dominated objects is clothing and accessories. Adorned personal (^$\hat{\mathcal{R}}_{f}\approx 77\%$) contains different types of jewelry, such as “crystal” or “ring”. Furniture (^$\hat{\mathcal{R}}_{f}\approx 64\%$) and textile (^$\hat{\mathcal{R}}_{f}\approx 62\%$) also fit this category, containing soft and textile-related objects such as “pillow” or “silk”. “Intimate wear” (^$\hat{\mathcal{R}}_{f}\approx 83\%$) contains underwear and swimwear. An additional female-leaning cluster, particularly in SD models, is “fruit” (^$\hat{\mathcal{R}}_{f}\approx 63\%$).
Most male-biased objects. Common male-dominated objects are audio speakers (audio, ^$\hat{\mathcal{R}}_{f}\approx 28\%$) and music instruments (music, ^$\hat{\mathcal{R}}_{f}\approx 27\%$), vehicles (transportation, ^$\hat{\mathcal{R}}_{f}\approx 26\%$), and metal objects (metal, ^$\hat{\mathcal{R}}_{f}\approx 21\%$). We see a clear contrast between female-dominated objects, which are fashion-related, and male-dominated objects, which are technical. The male dominance in musical instruments is unexpected, as they oppose existing gendered associations of certain musical instruments [4, 3]. While we see these gender associations reflected in higher female ratios relative to other instruments (see Sec. E.2), musical instruments remain male-dominated.
As shown in Fig. 3, occupation prompts lead to extreme ratios of generated female- or male-gendered images. There are comparatively few female-dominated occupations, while most occupations are strongly male-dominated.
Most female-dominated occupations. We find many relations to previously discussed female-dominated prompt clusters. For example, veterinary jobs (^$\hat{\mathcal{R}}_{f}\approx 90\%$) echo the pet-care-related activities, which we show are strongly female-dominated in Sec. 4.1. Nursing jobs (^$\hat{\mathcal{R}}_{f}\approx 96\%$), care jobs (^$\hat{\mathcal{R}}_{f}\approx 86\%$), and social jobs (^$\hat{\mathcal{R}}_{f}\approx 90\%$) all describe human-centered caring activities. Dental jobs (^$\hat{\mathcal{R}}_{f}\approx 84\%$) contains 4 occupations: “dental hygienist” and “dental assistant” are strongly female-dominated across all T2I models. “Dental technician” and “dentist” are female-dominated for all models except Flux-Schnell and SD-3.5-Large (see Sec. E.3 for detailed values). In the U.S., “dental hygienist” and “dental assistant” are occupations where $>90\%$ of the workforce are women, whereas $\approx 59\%$ of dental technicians and $\approx 40\%$ of dentists are women [64]. In Sec. F.3, we discuss bias amplification in T2I models regarding occupations by analyzing how gender ratios in images generated by models compare to actual workforce representation. Specifically, in Tab. 7, we show that T2I models not only replicate but often amplify existing gender skews found in occupations.
Most male-dominated occupations. While many occupations are male-dominated, the most male-dominated occupations are blue-collar jobs involving physical labor. This is, for example, the case with installer jobs (^$\hat{\mathcal{R}}_{f}\approx 0\%$), construction jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$), and wood jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$). Wood jobs comprises occupations such as “carpenter” and “sawing machine operator”. Similarly, vehicle jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$) and rail jobs (^$\hat{\mathcal{R}}_{f}\approx 1\%$) are transportation industry occupations. In contrast to female-dominated occupations, none of the top male-dominated occupations are human-centric occupations.
4.3 Special Topics
We now look closely at specific topics within the activity prompt group that are particularly relevant to societal impact, namely household and work/money-related activities.
Household Chores.
Figure 6: (Top) Household-related clusters of activity prompts. (Bottom) Work and money-making related clusters of activity prompts.
The division of household chores between spouses in heterosexual marriages is strongly moderated by gender [22, 44, 14, 51] and is relatively constant over time [26]. To select a subset of household-chores-related clusters, we classify all prompts in the activities prompt group as representing a household chore or not by an LLM (see Sec. F.1), specifically Phi-4. We cluster the resulting 105 activities and get 14 clusters, which we label manually and plot in Fig. 6 (Top).
The seven clearly female-leaning clusters are laundry (^$\hat{\mathcal{R}}_{f}\approx 68\%$), dishes (^$\hat{\mathcal{R}}_{f}\approx 66\%$), making bed (^$\hat{\mathcal{R}}_{f}\approx 65\%$), clean bathroom (^$\hat{\mathcal{R}}_{f}\approx 62\%$), organizing (^$\hat{\mathcal{R}}_{f}\approx 57\%$), cleaning/vacuuming (^$\hat{\mathcal{R}}_{f}\approx 56\%$), and sweeping floor (^$\hat{\mathcal{R}}_{f}\approx 56\%$). Note that models are not uniformly biased in the categories, but generally, SD-3.5-Large and Flux-Schnell exhibit fewer biases than other T2I models in that there are more men in images representing these tasks. All these clusters are related to various forms of cleaning. If we compare with the typical household chore division [22], we find that most female-typed and shared cleaning chores are female-dominated in T2I models, e.g., “making bed” and “vacuuming” are listed as shared chores in [22], but “making bed” is female-dominated in images from by Flux variants. “Vacuuming” is female-dominated in SD variants.
On the other end of the spectrum, we find the male-dominated clusters outside work (^$\hat{\mathcal{R}}_{f}\approx 5\%$), containing activities such as “working on the house”, and mowing lawn (^$\hat{\mathcal{R}}_{f}\approx 15\%$). Both are also more frequently performed by men [22]. This is also true for trash/cleaning (^$\hat{\mathcal{R}}_{f}\approx 42\%$), which, however, in our case, is a heterogenous cluster and also contains “doing a ton of spring cleaning” and “doing daily housework”, which are female-dominated in T2I models, while trash-related activities are strongly male-dominated. Interestingly, watering lawn (^$\hat{\mathcal{R}}_{f}\approx 46\%$) is male-typed in [22], but not clearly male-dominated in T2I models. Other not strongly gender-associated clusters are also listed as shared in [22].
Work and Money-Making. To assess gender bias regarding work or money-making-related activities, we also classify all 1405 activities by Phi-4 (see Sec. F.1) and cluster the resulting 139 prompts. This results in 20 clusters, which we label manually and show in Fig. 6 (Bottom).
No cluster other than teaching (^$\hat{\mathcal{R}}_{f}\approx 63\%$), work with animals (^$\hat{\mathcal{R}}_{f}\approx 62\%$), and writing (^$\hat{\mathcal{R}}_{f}\approx 59\%$) contains a majority of female-dominated activities. The cluster with the highest ratio of female-gendered images is teaching, which reflects our previous finding that teachers are associated with women. As already seen in Sec. 4.1, pet-related activities are frequently associated with women, and linking women with writing resembles the finding that humans associate women more than men with arts [63, 13]. Most other money-making activities, including regular work (^$\hat{\mathcal{R}}_{f}\approx 34\%$) and money-making (^$\hat{\mathcal{R}}_{f}\approx 24\%$), which refers to general activities related to money such as “worrying about money and time”, are male-dominated. We only see higher female ratios for job-seeking activities, i.e. job application (^$\hat{\mathcal{R}}_{f}\approx 46\%$) and job interview (^$\hat{\mathcal{R}}_{f}\approx 41\%$). This is concerning as underrepresenting women in work- and business-related contexts could reinforce existing stereotypes about women’s role in the workforce, perpetuating or even amplifying limiting gender norms of women as caretakers and men as breadwinners.
5 Discussion and Limitations
Analyzing many prompts and corresponding concepts, a clear picture emerges of which activities, contexts, objects, and occupations T2I models associate most with women or men. First of all, T2I models tend to generate more images of men than women. This highlights how men are the general default in model outputs, including in gender-neutral prompts and categories. The scenarios where men are not the default are in prompts of activities, contexts, objects, and occupations that are stereotypically feminine. We consistently observe that scenarios with a high rate of female-gendered images portray women in traditional roles: as homemakers, while shopping, or engaged in arts and beauty in our activity prompts; as caring and service-oriented in our contexts and occupations; and with fashionable and soft objects. This contrasts with male-dominated scenarios. Men are associated with physical work, both in the household and at their jobs, working with machinery, and are strongly associated with business. While this reflects the greater numbers of women in caretaking roles and men in machinery-related or business leadership roles that exist in society, our analysis clearly shows that gender stereotypes are replicated, reinforced, and further amplified in T2I models. While we can only numerically prove bias amplification in occupations due to the existence of workforce labor statistics to compare against which do not exist for other scenarios (see Sec. F.3), we hypothesize that the same is happening for other scenarios beyond occupations.
Our study offers detailed and valuable insights into gender bias in T2I models, which should be revisited and highlighted repeatedly [2]. However, our focus on binary gender leaves room for further exploration of social biases in T2I models. Regarding validity, we take concrete steps to include only prompts and images suitable for evaluating gender bias through person detection and filtering. Although our analysis relies on automatic methods, the strength of the evidence supporting our conclusion leaves little doubt that the observed effects stem from spurious biases in automatic processing. This study does not examine how prompts that evoke particular gender identities (e.g., a female engineer) can solve some of these issues, or whether such outputs have other gendered stereotypes. Rather, our contribution is analyzing outputs for gender-neutral prompts to unpack underlying defaults and gender biases present in models.
6 Conclusion
We present the first large-scale analysis of gender bias in T2I models related to everyday activities and contexts. We generated 3,217,000 images (2,293,295 after filtering) for 3,217 prompts covering activities, contexts, objects, and occupations. Based on this extensive dataset, we conduct a detailed analysis of gender bias in the generated images. We find that models default to men in outputs, including for gender-neutral categories. Further, we find that models replicate and amplify gender norms and stereotypes; current models risk reinforcing and further solidifying harmful biases or limiting norms through widespread deployment. While models differ in the bias patterns they exhibit, likely also due to different training data, we do not observe that larger models are noticeably less biased than smaller or less capable models, suggesting that gender bias cannot be addressed by simply increasing the models’ capacity. Our research addresses a key gap in the literature on social bias in T2I models, significantly expanding and empirically grounding our understanding of gender stereotypes in synthetic images. Further research should incorporate intersectionality and explore how non-binary people are presented, including through prompts that explicitly evoke different gender identities and intersectional categories.
Acknowledgements
This work was partially funded by the ERC (853489 - DEXIM) and the Alfried Krupp von Bohlen und Halbach Foundation, which we thank for their generous support. The authors gratefully acknowledge the scientific support and resources of the AI service infrastructure LRZ AI Systems provided by the Leibniz Supercomputing Centre (LRZ) of the Bavarian Academy of Sciences and Humanities (BAdW), funded by Bayerisches Staatsministerium für Wissenschaft und Kunst (StMWK).
Supplementary Material
Appendix A Prompts
In Sec. 3.1, we describe how we process prompts used to generate images and how we generate prompt variations. Here, we give further details on the precise prompts we use, and we also compare different LLMs to process prompts, justifying our choice of Yi-1.5-34B.
A.1 Prompt Processing LLM Prompts
In the following, we list the prompts used to process activities, contexts, objects, and occupations. All prompts are processed by Yi-1.5-34B and we use the following system prompt:
You are a helpful assistant that writes short sentences.
Activities. We use the following prompt for our LLM to process activities:
Examples:
- a person is sleep -> a person is sleeping
- a person is clean the house -> a person is cleaning the house
- a person is call dad -> a person is calling dad
- a person is craft ones inspiration angels -> a person is crafting their inspiration angels
- a person is kiss ones spouse -> a person is kissing their spouse
Rewrite this following the examples:
a person is {activity}’ ->
Note that line breaks are inserted automatically. The goal is mainly to generate syntactically correct prompts by properly inflecting verbs and changing word order and pronouns accordingly. “{activity}” is replaced by the respective activity phrase. We provide few-shot examples to guide the LLM.
Contexts. We use the following prompt for our LLM to process contexts:
Examples:
- a person <PREP> the alley -> a person in the alley
- a person <PREP> the wind farm -> a person near the wind farm
- a person <PREP> the piano story -> a person inside the piano story
- a person <PREP> the church -> a person in front of the church
- a person <PREP> the hospital -> a person at the hospital
Rewrite this following the examples:
a person <PREP> the {context}’ ->
The goal is to insert prepositions that match the given context. “{context}” is replaced by the respective given context from the SUN database. We provide few-shot examples to guide the LLM.
Objects. We use the following prompt for our LLM to process objects:
Examples:
- a person and a skis -> a person and skis
- a person and a airplane -> a person and an airplane
- a person and a sports ball -> a person and a sports ball
Rewrite this following the examples:
a person and a {object}’ ->
The goal is to insert the correct article for the given object. “{object}” is replaced by the respective given object. We provide few-shot examples to guide the LLM.
Occupations. We use the following prompt for our LLM to process occupations:
Examples:
- Management occupations -> manager
- Miscellaneous health technologists and technicians -> health technologist
- Animal control workers -> animal control worker
- Embalmers, crematory operators, and funeral attendants -> funeral attendant
- Sales representatives, wholesale and manufacturing -> sales representative
- First-line supervisors of construction trades and extraction workers -> construction supervisor
- Carpet, floor, and tile installers and finishers -> carpet installer
- Other healthcare practitioners and technical occupations -> healthcare practitioner
- Sales and related workers, all other -> sales representative
Summarize this occupation following the examples:
’{occupation}’ ->
The goal is to summarize and simplify lengthy occupation descriptions from the U.S. Bureau of Labor Statistics occupation list. “{occupation}” is replaced by the corresponding given occupation. We provide few-shot examples to guide the LLM. The generated occupation summary is inserted into the following template:
a person working as {occupation}
A.2 Prompt Processing LLM Comparison
Rewriting our prompts only requires shallow syntactical rewriting; therefore, we do not require particular reasoning skills from the LLM. Since we provide few-shot examples, we think most LLMs are suitable for our prompt processing. We decided to use Yi-1.5-34B due to its satisfactory performance. However, we compared four popular LLMs on 5 randomly sampled activities and found that all yielded the same results. In all cases, we used exactly the same prompts and system prompts. The compared LLMs are Yi-1.5-34B [98], Qwen2.5-32B-Instruct [97], Llama-3.1-8B-Instruct [27], and Phi-4 [1]. The results are in Tab. 3. All row-wise entries are identical, except for “drive around to look at sights with family in new home area” doesn’t insert the pronoun “their” before “family” in the processed prompt. We conclude that the choice of LLM is not crucial for our purposes, and we do not expect significant differences when using a different LLM than Yi-1.5-34B.
Original Yi-1.5-34B Qwen2.5-32B Llama-3.1-8B Phi-4
watch documentaries a person is watching documentaries a person is watching documentaries a person is watching documentaries a person is watching documentaries
drive around to look at sights with family in new home area a person is driving around to look at sights with their family in the new home area a person is driving around to look at sights with their family in the new home area a person is driving around to look at sights with family in the new home area a person is driving around to look at sights with their family in the new home area
brush one’s teeth a person is brushing their teeth a person is brushing their teeth a person is brushing their teeth a person is brushing their teeth
go to the pet blessing at church a person is going to the pet blessing at church a person is going to the pet blessing at church a person is going to the pet blessing at church a person is going to the pet blessing at church
go to get lunch and froyo with a friend on the weekend a person is going to get lunch and froyo with a friend on the weekend a person is going to get lunch and froyo with a friend on the weekend a person is going to get lunch and froyo with a friend on the weekend a person is going to get lunch and froyo with a friend on the weekend
Table 3: Comparison of 4 different LLMs (Yi-1.5-34B [98], Qwen2.5-32B-Instruct [97], Llama-3.1-8B-Instruct [27], and Phi-4 [1]) on 5 randomly sampled activities. Prompts and system prompts are the same in all cases. All LLMs lead to the same processed prompts, suggesting the choice of LLM is irrelevant for our prompt processing purposes.
A.3 Prompt Clustering
HDBSCAN Settings. As described in Sec. 3.1, we use the HDBSCAN clustering algorithm to cluster prompt embeddings. Prompt embeddings are obtained from the all-mpnet-base-v2 model provided by [70] and reduced to 16 dimensions by UMAP [62]. For HDBSCAN, we use the implementation from scikit-learn with the following parameters:
min_cluster_size 3
min_samples 3
metric cosine
cluster_selection_method leaf
All other parameters are the default parameters of the scikit-learn implementation. The parameters have been manually selected, which leads to a very fine-grained clustering, which is intended.
Cluster summarization. We summarize prompt clusters using an LLM. An example of summarization is in Tab. 5. Concretely, we use Llama-3.3-70B-Instruct [27], with the following prompt:
Consider the following {prompt_group}:
{prompts}
Give a short and descriptive title of the complete list. When creating the title, follow these guidelines:
- Capture the essence of the whole list, not individual {prompt_group}.
- Ensure the title accurately reflects all the {prompt_group} in the list.
- Keep it concise, using 3 words or fewer.
- Do not add information that is not present in the list.
- Avoid adjectives or qualifiers that are not explicitly mentioned.
- Be as precise as possible and avoid being overly general.
- The title should end with {specifier}.
Your summary:
The placeholder {prompts} is replaced by the list of prompts that we want to summarize, and each prompt appears in a new line. The values of {prompt_group} and {specifier} are taken from the following table, which maps prompt groups to the respective values:
Prompt Group {prompt_group} {specifier}
Activities activities activities
Contexts contexts places
Objects objects objects
Occupations occupations jobs
We decide to use Llama-3.3-70B-Instruct after comparing to other state-of-the-art LLMs, namely Qwen2.5-72B-Instruct [97], Yi-1.5-34B [98], and Phi-4 [1]. A comparison of the LLMs on 4 illustrative samples (activity clusters) is in Tab. 4. We notice that Llama-3.3-70B is superior in terms of how concise and fluent the resulting summaries are.
Id. Llama-3.3-70B Qwen2.5-72B Yi-1.5-34B Phi-4
21 Email activities Emailing for activities Email Correspondence Activities Email Writing Activities
46 Work activities Work and Meetings Activities Professional Engagement Activities Professional and Social Activities
96 Baking activities Baking Sweet Activities Baking and Dessert Activities Baking and Baking Activities
144 Reading activities Diverse Reading Activities Versatile Reading List Diverse Reading Activities
Table 4: Comparison of cluster summaries generated by different LLMs. Summaries generated by Llama-3.3-70B stand out for being both concise and linguistically fluent.
Prompts Summary
shopping at walmart Grocery shopping
doing grocery shopping
going grocery shopping
shopping for groceries
going shopping for groceries
going shopping at the grocery store
Table 5: Example cluster and cluster summary. On the left, we show the prompts in the cluster, omitting the prefix “a person is”.
Appendix B Image Generation
B.1 Diffusion Model Settings
Generally, we use the hyperparameters (guidance scale, number of diffusion steps) recommended by the model authors. In all cases, the number of diffusion steps is 50, except for Flux-Schnell, where being a few-step-model [76] enables generating images with 4 diffusion steps. Guidance scales are as follows: 3.5 (Flux); 0.0 (Flux-Schnell); 3.5 (SD-3.5-Large); 4.5 (SD-3.5-Medium); and 7.0 (SD-3-Medium). Images are generated in $1024\times 1024$ for all models except Flux, where we generate images in $512\times 512$ for improved generation efficiency. After generation, all images are downscaled to $512\times 512$. Also note that, for Stable Diffusion models, we add the prompt prefix “a high-quality picture of” as we found this improves generation quality.
B.2 Prompt Following
We use VQAScore [54] to measure how if generated images match their respective given prompts. VQAScore has been shown to yield better performance than related measures such as CLIPScore [41] or TIFA [47]. VQAScore uses an MLLM (clip-flant5-xxl which was trained by the authors of VQAScore specifically for this purpose) to predict the probability of answering “yes” when providing the MLLM the image and the following prompt:
Does this figure show "{prompt}"? Please answer yes or no.
where “{prompt}” is replaced with the actual prompt used to generate the image. This yields a probability between 0 and 1, where a higher value indicates a stronger agreement between the prompt and the image. Therefore, higher values of VQAScore are desirable when generating images. However, models are expected to not generate good images for all prompts, and VQAScore is based on a statistical model with its own failure modes, introducing error-compounding effects. Also, for these purposes, it is not strictly necessary that models generate images that faithfully depict the prompt. We are interested in the associations of T2I models and gender, not general image quality. In Fig. 7, we show summary statistics of VQAScore values factored by combinations of models and prompt groups. We can see that in most cases, the VQAScore is above 0.5, indicating good prompt following.
Figure 7: Distributions of VQAScore values [54] factored by combinations of models and prompt groups. Higher scores are better.
B.3 Effect of Prompt Variations
In Fig. 8, we study the effect of our 5 different prompt variations on gender. Concretely, we calculate the deviation from the mean female ratio across all variations for each individual prompt variation. Then, we plot the resulting values factorized by T2I model. We see that variations have slight individual effects on the gender, but they are balanced. No prompt variation significantly skews the gender distribution towards one gender across all prompts. The strongest effects are observed for the “individual” variation, which leans more toward men than other variations. “person” is leaning more toward female-gendered images than the average. Overall, we conclude that the validity of our results is not affected by our different prompt variations.
Figure 8: Effect of different prompt variations (“a person”, “an individual”, …) on the ratio of female-gendered images, factorized by models. Positive values indicate female skew, while negative values indicate male skew compared to the average across variations.
Appendix C Gender Indentification
C.1 MLLM Prompt
To identify perceived gender, we use the InternVL2-8B model. The InternVL2 model series [17, 18] was the strongest open model series when conducting experiments. We chose the 8B variant as it offers the best performance-efficiency tradeoff. Larger models do not perform better at perceived gender classification but incur a significant computational overhead.
We use the following prompt to identify gender:
What is the gender of the person in the image?
A. female
B. male
C. unclear/cannot tell
Answer with a letter (A, B, C, etc.).
Additionally, we randomly permute the option order (but not the letter order) to avoid label bias (e.g. the model preferring to predict the option letter “A”) [25].
We validate the performance of InternVL2-8B on VisoGender [38]. All images were labeled by human annotators for perceived gender. Specifically, we predict the gender of all 229 images in VisoGender that show a single person. 228 predictions are correct, meaning that perceived gender can be identified nearly perfectly by InternVL2-8B.
C.2 Nonbinary Gender Labels
We also evaluate if InternVL2-8B labels images as “nonbinary”. For this, we repeat the gender identification described in Sec. C.1, but add “nonbinary” as an option in addition to “female”, “male”, and “unclear/cannot tell”. Among the 5,675,715 person bounding boxes, only 332 receive the label “nonbinary”. This is not enough to conduct a meaningful quantitative analysis. However, we show 19 of 20 unique images generated by Flux that receive the “nonbinary” label. We removed one image that shows NSFW content. The images are in Fig. 9.
Figure 9: Images generated by Flux that receive “nonbinary” as perceived gender.
C.3 Images without Recognizable Gender
In Sec. 3.3, we filter images that show people but no person has a clearly recognizable gender according to InternVL2-8B. In total, 302,829 images are filtered by this criterion. One concern is that the gender of people in these images is perceived as nonbinary. Therefore, we inspect a sample of the filtered images but find that they are images where no gender cues are visible due to occlusion (shade, clothes), small size of people, or blurriness of people (in the background). Other images show only body parts, infants, or nonhuman creatures. We display 10 examples in Fig. 10.
Figure 10: Example (Flux) images filtered because the shown people’s gender is uniformly labeled “unclear/cannot tell” by InternVL2-8B. Detected person bounding boxes are in red. Examples include small, blurry, or occluded people, as well as infants, body parts or nonhuman creatures. We do not find evidence of images showing nonbinary gender.
Appendix D Bias Agreement across Models
For each prompt group, we calculate the Spearman correlation between female ratios for all pairs of models. Correlations are only calculated on prompts that are not filtered for any model to ensure comparability. Results are in Fig. 11. We can see that all correlations are very high, especially for occupations and activities.
Figure 11: Spearman correlation of model pairs across female ratios in all prompt groups.
Appendix E Detailed Analyses of Clusters
E.1 Retail contexts
In Sec. 4.1, we observe that places in the “retail” cluster are partially strongly female-dominated. Here, we further prove that female-dominated places predominantly relate to fashion, clothes, and beauty. To this end, in Fig. 12, we plot all places in the “retail” cluster where at least one of the five T2I models generated 60% or more female-gendered images. There, we observe that of 14 places, 9 are related to fashion, beauty, or luxury (“beauty salon”, “sweing room”, “dress shop”, “perfume shop”, “wig shop”, “fitting room”, “jewelry shop”, “clothing store”, “fabric store”). In particular, this comprises the most female-dominated retail places.
Further retail places include shopping-related (“drugstore”, “department store”), which we identified as female-associated activity in Sec. 4.1, and “florist shop”, which relates to flowers being a female-leaning type of object.
Figure 12: Detailed breakdown of places in the retail cluster.
E.2 Music Instruments
In Sec. 4.2, we find that music instruments make up a male-dominated cluster. This is surprising, as previous research found clear gender associations with respect to musical instruments. In Fig. 13, we show the ratios of female instruments for all objects in the “music instruments” cluster. Note that the objects disc and slipknot are not musical instruments, but we show them nonetheless because they are included in the cluster. The relatively most female-leaning instruments are flute and violin, in accordance with [4, 3]. The same is true for drum, saxophone and guitar, which are male-leaning. However, as also noted in Fig. 5, overall musical instruments are male-leaning and do not follow the associations made by humans.
Figure 13: Detailed breakdown of gender ratios of objects in the music instruments cluster.
E.3 Dental Jobs
In Sec. 4.2, we take a closer look at the four occupations clustered as “dental jobs”. In 3 of the 4 occupations, the majority of the workforce in the U.S. is women ($>90\%$ for dental hygienist and dental assistant, and $\approx 60\%$ for dental technician) [64]. $\approx 40\%$ of dentists are women. These patterns are reflected in the ratios of female-gendered images generated by T2I models, as shown in Fig. 14. However, SD-3.5-Medium and SD-3-Medium are significantly more biased towards generating female-gendered images than other models.
Figure 14: Detailed breakdown of gender ratios of occupations in the dental jobs cluster.
Appendix F Special Topics
F.1 Activity Classification
For our analyses in Sec. 4.3, we classify our 1405 activities by an LLM to determine if they relate to household chores and work/money. To conduct the classification, we compare 4 different LLMs, namely Llama-3.3-70B-Instruct [27], Qwen2.5-72B-Instruct [97], Yi-1.5-34B [98], and Phi-4 [1]. For classification into household chores, we use the following prompt:
Is the following activity considered a household chore: {activity}. Answer yes or no.
and for classification into work/money-related actviities we use
Is the following activity related to paid work or money-making (not household work, shopping, or hobbies): {activity}. Answer yes or no.
In both cases, we replace {activity} with the activity prompt that is to be classified. Also, we always use the following system prompt:
You are a helpful assistant that writes short sentences.
In Tab. 6, we show the number of activities that are classified as being related to the two categories, i.e. where the model answers “yes”. Phi-4 labels the fewest activities as household-related or work/money-related, and thus, we proceed with this model, as a lower number of activities makes the analysis more comprehensive. A manual analysis also suggests that the precision of Phi-4 is better than the precision of other models.
Llama-3.3-70B Qwen2.5-72B Yi-1.5-34B Phi-4
Household 203 155 205 105
Work/Money 245 192 187 148
Table 6: Number of activities (out of all 1405 activities) classified as household chores or work/money-related by different LLMs. Phi-4 yields the fewest activities in both categories.
F.2 Work-related contexts
We further analyze work-related places in the contexts prompt group. To select work-related places, we classify all 737 contexts by 4 LLMs (Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct, Yi-1.5-34B, and Phi-4) and continue to work with the classifications from Yi-1.5-34B, which yields the best precision upon manual inspection. The prompt used to obtain labels for context is
Is the following place related to paid work or money-making (not household work, shopping, or hobbies): {place}. Answer yes or no.
where we replace {place} with the context to be classified. We then cluster the resulting 156 work-related places with the method described in Sec. 3.1 and obtain 24 clusters. These are shown in Fig. 15 together with the respective per-model female ratios.
We notice that most clusters are male-dominated, in line with our findings in Sec. 4.1 and Sec. 4.2. The male dominance is particularly strong in clusters related to transportation (shipping) and industrial sites (factories, power/gas/recycling, mine/excavation, …). Female ratios are comparatively higher in contexts related to art (art/television), shopping (market/shop, fashion shops), and places of pleasure (hotel/casino). This confirms our observation that T2I models reflect gender stereotypes and associated work, especially physical labor, with men and social places with women.
Figure 15: Female ratios of clusters obtained for places classified as work-related. Error bars refer to the standard deviation of female ratios across all prompts contained in the respective cluster.
To narrow down the analysis to office-related places, which are subsumed together with many unrelated places in the misc (office cluster in Fig. 15, we further classify contexts as office-related by Yi-1.5-34B using the following prompt:
Is the following place related to office work, meetings, or conferences: {context}. Answer yes or no.
We show the resulting 14 places alongside the per-model female ratios in Fig. 16. There, we find that most office-related places are male-dominated. Generally, SD models have higher female ratios across all places. Places with comparatively high female ratios are “call center” and “reception”, which are related to professions where the majority of the workforce are women: Call center employees are listed under “Customer Service Representatives” by [64], and the ratio of women in the U.S. is $65.2\%$. Also, 89.1% of receptionists are women. In SD models, “breakroom” and “office cubicles” are also gender-balanced. In conclusion, the closer analysis of office-related places further strengthens the impression that T2I models associate work more with men than with women.
Figure 16: Ratios of places classified as office-related.
F.3 Bias Amplification in Occupations
Here, we analyze the relationship between gender ratios in images generated by T2I models and the actual representation of women in the U.S. workforce, as reported by [64]. Of the 575 occupations in our study, [64] provides the percentage of women for 365 occupations. For each occupation prompt $p$, we compute
$\Delta(p)=\mathcal{R}_{f}^{\textrm{bls}}(p)-\mathcal{R}_{f}(p)\in[-1,1]$ (4)
where $\mathcal{R}_{f}^{\textrm{bls}}(p)$ represents the proportion of women in the U.S. workforce. A positive $\Delta$ indicates that T2I models generate fewer women than the actual workforce proportion, while a negative $\Delta$ indicates that they generate more women than expected. In Fig. 17, we present the distributions of $\Delta$ values for all five T2I models. Overall, the distributions tend to be centered above zero, indicating that, on average, T2I models depict a higher proportion of men compared to actual workforce statistics.
Figure 17: Distribution of differences $\Delta$ between female ratios in occupation images generated by T2I models and real-world (U.S.) ratio of women in the workforce for the respective occupation. Positive values indicate more men in generated images than in the workforce, and negative values indicate more women in generated images than in the workforce.
To further explore this perspective, we analyze bias amplification in occupations based on whether the majority of the workforce is male or female. This analysis is presented in Tab. 7. First, we classify each occupation as either “male majority” or “female majority” based on the actual proportion of women in that occupation. If more than 50% of the workforce is female, the occupation is labeled as “female majority”; otherwise, it is labeled as “male majority”. Next, we examine whether the proportion of men or women increases or decreases in images generated by T2I models. If the ratio of the majority group increases, we say the bias is amplified, whereas if it decreases, we say the bias is reduced.
From Tab. 7, we observe that bias in male-majority occupations is almost always amplified. For Flux models, bias in female-majority occupations is more often amplified than reduced. In contrast, for Stable Diffusion models, bias in female-majority occupations is more often reduced than amplified. Overall, these findings confirm our observation that T2I models tend to increase the proportion of men in generated images, while also showing numerous cases where female-majority bias is amplified.
Male majority Female majority
reduced amplified reduced amplified
Flux 7.12% 44.11% 19.18% 29.59%
Flux-Schnell 6.03% 45.21% 17.81% 30.96%
SD-3.5-Large 7.67% 43.56% 23.01% 25.75%
SD-3.5-Medium 10.68% 40.55% 28.22% 20.55%
SD-3-Medium 9.32% 41.92% 28.77% 20.00%
Table 7: We classify occupations into “male majority” or “female majority” based on whether there are more men than women in the workforce (actual U.S. statistics). Then, we check if, in generated images, the majority gender has increased or decreased ratio. If the majority gender is increased, we label it as “amplified”; if the majority gender is decreased, we label the occupation as “reduced”.